OpenStudy (anonymous):

The electric force between a charged foam cup and a positively charged metal sphere is 2.12 newtons. If the charge on the cup is 2.0 × 10-6 coulombs and the charge on the metal sphere is 2.5 × 10-6 coulombs, what is the distance between the two? (k = 9.0 × 109 newton·meter2/coulombs2) i know the formula but dont know how to use it :(

OpenStudy (johnweldon1993):

Hint, Coulombs Law states \(\large F = k\frac{|q_1||q_2|}{r^2}\) Where F is the force between the 2 objects (given) Soo we can pug everything we have in \[\large 2.12N = (9.0\times 10^9 \frac{Nm^2}{C^2})(\frac{|2.0 \times 10^{-6}C||2.5\times 10^{-6}C|}{r^2})\] So notice that we just want to solve for the distance between the 2...hence we just solve for 'r' \[\large r = \sqrt{\frac{ (9.0\times 10^9 \frac{Nm^2}{C^2})|2.0 \times 10^{-6}C||2.5\times 10^{-6}C|}{2.12N}} = ?\]